OpenStudy (anonymous):

A plate has a diameter of 14 inches. What is the approximate circumference of the plate? A. 21.98 in. B. 43.96 in. C. 87.92 in. D. 153.86 in. PLEASE HELP ASAP

OpenStudy (mathstudent55):

The circumference of a circle is \(C = \pi d\) where C = circumference \(\pi \approx 3.14 \) d = diameter

OpenStudy (anonymous):

why is the plate a foot long? is the real question

OpenStudy (mathstudent55):

\(14~in. \ne 1~ft\) \(14~in. \gt 1~ft\) Two possibilities: 1. A big eater. 2. A serving plate.

OpenStudy (anonymous):

these answers what

OpenStudy (mathstudent55):

\(C = \pi d\) \(C = 3.14 \times 14~in.\) Multiply the numbers together.